Convert the following numbers to hexadecimal (base 16). Show your work.

1   1110101012  base 2

2   1011012 base 2

3 5510 base 10

4 110 base 10

5 5210 base 10
Answer #1
Convert each 4 bits of binary to its hexa decimal

Binary to hexa:
--------------------
0000 in hexa is 0
0001 in hexa is 1
0010 in hexa is 2
0011 in hexa is 3
0100 in hexa is 4
0101 in hexa is 5
0110 in hexa is 6
0111 in hexa is 7
1000 in hexa is 8
1001 in hexa is 9
1010 in hexa is A
1011 in hexa is B
1100 in hexa is C
1101 in hexa is D
1110 in hexa is E
1111 in hexa is F

============================

1   1110101012  base 2
convert each 4 bits to hexa.
1101 in hexa is D
0101 in hexa is 5
So, Hexa for the binary 1 1101 0101 is 1D5

2   1011012 base 2
convert each 4 bits to hexa.
1101 in hexa is D
So, Hexa for the binary 10 1101 is 2D

3 5510 base 10
55/16 = 3 and reminder is 7
3/16 = 0 and reminder is 3
so, 55 in hexa is 37

4 110 base 10
1 in hexa is 1.

5 5210 base 10
52/16 = 3 and reminder is 4
3/16 = 0 and reminder is 3
so, 52 in hexa is 34